OpenStudy (anonymous):

Solve the following system of linear equations. 3x + 2y = 10 2x + 3y = 15/2

OpenStudy (anonymous):

By Gaussian reduction: \[ \begin{align} 3(3x+2y)&=2(10)\\ 2(2x+3y)&=2(\frac{15}{2}) \end{align} \]Which becomes: \[ \begin{align} 9x+6y&=30\\ 4x+6y&=15 \end{align} \]Subtracting one from the other gives us one nice equation: \[ 5x=15\implies\\ x=3 \]Plugging this value into the above: \[ 3(3)+2y=10\implies\\ 2y=1\implies\\ y=\frac{1}{2} \]As a check: \[ 2(3)+3(\frac{1}{2})=6+\frac{3}{2}=\frac{15}{2} \]Hence, we are done.